A spaceship of mass m has velocity v in the positive x direction of an inertial reference

frame. A mass dm is fired out the rear of the ship with constant exhaust velocity (-v0) with

respect to the spaceship. a) using conservation of momentum, show that dv/v0 = dm/m, b)

By integration, find the dependence of v on m if v1 and m1 are the initial values. c) Can the

acceleration be constant if dm/dt, the burning rate is constant.


1
Expert's answer
2021-01-25T14:05:20-0500

Given,

The mass of the spaceship be m and velocity be v.

a) So, initial momentum of the spaceship be "(p)=mv"

Now, as the dm mass fired from the spaceship, then also momentum always be conserve

"\frac{dp}{dt}=0"


"\Rightarrow \frac{d}{dt}(mv)=0"


"\Rightarrow \frac{dm}{dt}v+m\frac{dv}{dt}=0"


"\Rightarrow \frac{dm}{dt}v=-\frac{dv}{dt}m"


"\frac{dm}{m}=-\frac{dv}{v}"


b) "\frac{dm}{dt}v+m\frac{dv}{dt}"

but here gravity is also working in downward direction,

"\int_{m_{1}}^{m-dm} \frac{dm}{m}=\int_{v_1}^v \frac{dv}{v}"


"\Rightarrow [\ln m]_{m_1}^{m_1-dm}=[\ln{v}]_{v_1}^v"

"\Rightarrow \ln[m_1-dm]-\ln[m_1]=\ln[v]-\ln[v_1]"

"\Rightarrow \frac{m_1-dm}{m_1}=\frac{v}{v_1}"

c)"m_1-dm=\frac{v}{v_1}m_1"

"dm=m_1-\frac{vm_1}{v_1}"

if dm/dt = constant

then "\dfrac{dv}{dt}=-g+k\frac{v'}{m}"